CRY+VIS Interface Meeting
Date: 2022/2/25 13:00-13:30
Participants: Takahashi,Uchiyama, Suzuki, Aso, Yokozawa, Yasui, Ikeda, Sato, Ushiba, Node, Akutsu
- IX: Glitch in the F3 signals. Windshield for the MN oplev. Cooling of the cryo-shield.
- EX: Health check.
- IY: Recovery of the IP.
- EY: Preparation for the accelerometers.
- Replacement of flip cables: Stock of the feedthrough flange.
- TM height setting: ITMY.
- Oplev: Setup of the Oplev in EY was in 70%. Not yet in IY.
- ACC: Installation was scheduled in the week of 2/28 in EY.
- Commissioning in air: Alignment of ITMY.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check: Cabling check
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Scheduled from 2/28 in EYV.
- Heater control: Test in IXV/EXV.
- Power source for Oplev: Check of ports is scheduled on 3/15.
- TMS: Installation of the balancer is scheduled in 2/28-3/2.
Date: 2022/2/25 14:00-14:30
Participants: Takahashi, Fabian, Ushiba, Uchiyama, Ishizaki, Aso, Tanaka, Sato
- Removed the ballast rings (2.7kg x2) in IY. The resonant frequencies went to 50mHz for L and 80mHz for T.
- Investigating the strange behavior of the IP. In the case of the working position deviating from the center, the lowest mode due to the filter chain disappeared.
- For the BS suspension, I calculated the decay time produced by the damper ring of the ~50 mHz yaw mode of the whole suspension: it is 31 s. Therefore, it's not necessary to adjust the position of the damper ring.
- I provided Marc with a Matlab version of the SR2 rigid body model. Back in the day I adjusted the parameters of the model for the calculated transfer functions to match the measured ones. Then I used it to calculate the noise budget that was shown in the Type-B paper.
- Last week, we recieved the cable clamps prototype for the uppermost statge of Type-Bp suspensions. We should test them very soon.
- Hirata-san designed two ways of mounting the limit switches in the traverser. After email discussion with Takahashi-san, Satou-san and me, it seems we'll make prototypes for both version to try.
In PRM, we examined more carefully the condition of H1 channel, whose actuation force seems unusually lower than in other channels. By exchanging cables with other coils, actuating on them sepatrately, and using the oplev TILT_VER channel, we realized that the problem is not within the chamber and it's likely in the coil driver. The effect of H1 on the mirror is at least 50% less of H1 compared with the effect of other coils.
Yokozawa-san reported that, in PRM, the length QPD board broke. Does anyone know the current status?
Type-B and Type-Bp TO-DO list
Test Type-Bp cable clamps prototype. This might be a messy task because some cables are entangled inside the chamber and we need to disentangle them. Which suspension should we use?
- Check the cable of the BS F0 fishing rod. There might be something wrong with it.
- I should provide diaggui templates for Type-B automeasurement. Currently, some from Type-A are in place for that.
- If possible, we should visually inspect the BS payload more carefully in case it's close to touching the security structure. After moving it in yaw for green laser alignment, Hirata-san only did a cursory check.
Not for VIS but we were assigned to do it: install the driver for the QPD in POP table.
- Type-A team: Investigation of GAS filters in IY. Installation of ACCs in EY.
- Type-B team: See TO-DO list.
(Date in Kamioka)
The week from 2/28
- Takahashi: 3/1-4
- Sato: 3/1-4
- Hirata: 3/1-4
The week from 3/7
- Takahashi: 3/8-11
- Sato: 3/8-11
- Hirata: 3/8-11